Find the length of the diagonal of a retangle with 5 inch width and 7 inch length. (rounding answer to thousandths
Answer by nyc_function(2733)

Use the Pythagorean Theorem: a^2 + b^2 = c^2.

Let a = 5 and b = 7.

5^2 + 7^2 = c^2

Let c = hypotenuse

Solve for c.

Can you do it now?